What is the heat of vaporization of a liquid that has a vapor pressure of 641 torr at 85.2
oC and a boiling point of 95.6 oC at 1.00 atm?
1
Expert's answer
2020-05-20T13:45:11-0400

"ln(P_2/P_1)=-\Delta H_v(1/T_2-1/T_1)/R"

"P_1 =641 torr=641atm\/760=0.8434atm"

"P_2=1atm"

"T_1=85.2^oC=358.35K"

"T_2=95.6^oC=368.75K"

"\Delta H_v=R*ln(P_1/P_2)/(1/T_2-1/T_1)="

"=8.31J\/(mol*K)*ln(0.8434)\/(1\/368.75K-1\/358.35K)="

"=17991.5J\/mol=17.991kJ\/mol"

"\Delta H_v=17.991kJ/mol"